OpenStudy (anonymous):

Which is a prime number? A. 15 B. 27 C. 41 D. 57

OpenStudy (solomonzelman):

A prime number is a number that you can't divide by any other number besides the number itself and 1.

OpenStudy (solomonzelman):

HINT: \(\Large\color{red}{ \bf 15=5 \times 3 }\) \(\Large\color{red}{ \bf 27=9 \times 3 }\) \(\Large\color{red}{ \bf 57=19 \times 3 }\)

OpenStudy (anonymous):

so 41?

OpenStudy (solomonzelman):

YES